### v1

@@ 5.4 Reflectance Singularities
In comparison with SuperNormal, our approach has the advantage of taking into account not only normals, but also reflectance. However, this may sometimes turn into a drawback. Indeed, the actual reflectance values weight the optimization process, potentially leading to singularities e.g., in the presence of very dark materials. Let us demonstrate this under the Lambertian assumption, in the case of grey scale images ($q=1$). - For simplicity, let us consider an ideal sampling scenario where the weight $w(\cdot,f(\mathbf{x}_{k}(\cdot)))$ in (9) is a Dirac function $\delta_{t_{k}}(\cdot)$, i.e. it is null everywhere except at some $t_{k}$ such that $\mathbf{x}_{k}:=\mathbf{x}_{k}(t_{k})$ lies on surface $\mathcal{S}$. By combining (3) and (10), the volumetric rendering (9) then simplifies to $\tilde{\mathbf{v}}_{k}(\boldsymbol{\rho},f)=\mathsf{L}_{k}\,\nabla f(\mathbf{x}_{k})\,\boldsymbol{\rho}(\mathbf{x}_{k})^{\top}$. If $q=1$, introducing the scalar variable $R$ such that $\boldsymbol{\rho}(\cdot)=\mathbf{r}\,R(\cdot)$, the NVR loss (8) can be rewritten:
$\mathcal{L}_{\text{NVR}}^{p}(R,f)=\sum_{k=1}^{m}|\mathbf{r}_{k}|\left\|R(\mathbf{x}_{k})\mathsf{L}_{k}\nabla f(\mathbf{x}_{k})-\mathsf{L}_{k}\,\mathbf{n}_{k}\right\|_{p}^{p}.$ (13)
This rewriting emphasizes that reflectance acts as weighting factor in the loss function. Consequently, singular reflectance values may substantially influence the optimization of geometry. In particular, dark colours essentially leave the geometry locally unconstrained, which may result in slow, sub-optimal convergence and a loss of fine details (see Figure 6-(b)).
@@
$\mathbf{r}_{k}$ $:=\frac{1}{q}\left[{\mathbf{r}_{k}}^{\top},~{}\left(q-\left\|\mathbf{r}_{k}\right\|_{p}^{p}\right)^{1/p}\right]^{\top}\,\forall k,$ (14)
$\boldsymbol{\rho}(\cdot)$ $:=\frac{1}{q}\left[\boldsymbol{\rho}(\cdot)^{\top},~{}\left(q-\left\|\boldsymbol{\rho}(\cdot)\right\|_{p}^{p}\right)^{1/p}\right]^{\top}.$ (15)
It is easy to check that the norm of these vectors is equal to 1, whatever the values of $p$ and $q$. This reflectance embedding strategy prevents any unintended influence of singular reflectance values on geometry optimization. This is qualitatively illustrated in Figure 6 on a synthetic experiment, using the $L_{1}$-norm ($p=1$) and grey level data ($q=1)$. Therein, the dark regions are not faithfully reconstructed in the standard case, whereas reflectance embedding significantly improves the results. Further experiments evaluating the interest of this embedding will be conducted in Section 6.
@@ 5.5 From RNb-NeuS to RNb-NeuS2
One of the main challenges with our initial approach, RNb-NeuS [7] built upon NeuS [82], was the significant computation time required for a single reconstruction. The scene-dependent process indeed usually takes approximately 6-8 hours on a NVIDIA RTX 4090 GPU. To address this issue, we restructured our algorithm to follow the NeuS2 [85] architecture. This results in a solution that is 100$\times$ faster, reducing computation time to around 5 minutes.
NeuS builds on NeRF [60] by introducing volumetric rendering with a transfer function that converts an SDF into a density function. This allows direct optimization of scene geometry and extraction of a mesh via the SDF’s zero-level set. NeuS2 applies the same principles but replaces NeRF with Instant-NGP [62], which enables real-time rendering thanks to CUDA acceleration and the use of an optimizable hash grid for sampling during volumetric rendering. While CUDA contributes the most significant performance improvement, the hash grid further reduces computation time by focusing on surface-adjacent regions, enhancing detail preservation. NeuS2 retains these key optimizations while also including the SDF-based density representation from NeuS.
@@ 6.1.1 Evaluation Datasets
DiLiGenT-MV [51] is a benchmark of five real-world objects, some with complex reflectance and surface profiles. Each object is imaged from 20 calibrated viewpoints using the classical turntable MVPS acquisition setup [33], under 96 directional lighting conditions. Given the acquisition characteristics, the relatively low image resolution of $612\times 512$ pixels corresponds to approximately $0.4$ mm per pixel.
@@ 6.1.2 Photometric Stereo Methods
Among these methods, SDM-UniPS [37] (referred to as SDM in the result subsections) provides both normal and reflectance maps, enabling a complete evaluation of our pipeline. UniMS-PS [31] predicts slightly more accurate normals (see Table 1), yet no reflectance. It thus allows only for normal integration assessment, similar to SuperNormal [10]. To establish an upper bound on the achievable reconstruction quality, experiments were also conducted using ground truth normals. In the forthcoming experiments, the notation N: X (resp. N/R: X) indicates that the inputs are normals alone (resp. normals and reflectance), estimated with method X.
Since architecture constraints limit SDM-UniPS to use only 10 PS images, a sampling strategy was employed. Specifically, each $\mathbf{r}_{k}$ and $\mathbf{n}_{k}$ was computed as the median of the estimated reflectances and normals across $100$ random trials, each trial involving 10 randomly chosen images. UniMS-PS scaling better thanks to multi-scale enhancement, sampling was not necessary.
@@ 6.2 MVPS Results
Methods Bear Buddha Cow Pot2 Reading Mean
Park16 0.932 0.375 0.335 0.972 0.525 0.628
Li20 $\dagger$ 0.206 0.257 0.101 0.213 0.255 0.206
Kaya22 0.380 0.396 0.297 0.390 0.344 0.361
PS-NeRF 0.253 0.304 0.275 0.248 0.353 0.287
Kaya23 0.320 0.202 0.209 0.370 0.270 0.274
MVPSNet 0.288 0.272 0.244 0.296 0.243 0.269
SuperNormal (N: SDM) 0.177 0.206 0.186 0.141 0.222 0.186
NLP-MVPS 0.211 0.176 0.169 0.213 0.250 0.204
RNb-NeuS (N/R: SDM) 0.255 0.216 0.282 0.175 0.276 0.241
@@
Table 2: Chamfer distance (lower is better) averaged overall all vertices on DiLiGenT-MV dataset. Best results. Second best. Since $\dagger$ requires manual efforts [51], it is not ranked.
@@ 6.2.1 Results on DiLiGenT-MV
Table 2 presents the CD evaluation for the DiLiGenT-MV dataset. As a preliminary observation, let us remark the significant discrepancy (around $-13\%$) between our results and those from the conference paper (RNb-NeuS), using the SDM-UniPS normal and reflectance maps. This discrepancy is due to the change of norm in the NVR loss (8), from $p=1$ to $p=2$, which severely impacts the CD scores. As shown in the ablation study (Table 8), using the $\mathcal{L}_{\text{NVR}}^{1}$ loss yields scores comparable with the conference paper.
Another notable observation is that whatever the PS method, our results are on par with state-of-the-art, or even better. Nevertheless, combining our method with UniMS-PS yields better results ($-14\%$) than with SDM-UniPS, consistent with the higher accuracy of the former’s normals (see the first row of Table 1). This highlights the versatility of the proposed method, which adapts well to PS advancements.
On this dataset, reflectance seems not to really improve the results, yet its importance will become clearer later, when considering datasets exhibiting more challenging materials.
@@
6.2.2 Results on LUCES-MV
Experiments on LUCES-MV dataset (Table 4) confirmed our state-of-the-art results, yet here reflectance visibly improves the results. Interestingly, the CD scores largely increase in comparison with DiLiGenT-MV despite a better pixel resolution: for the Bowl object, even using GT normals, the CD is above $0.3$ mm. We believe this could be due both to the sparser number of views — 12 vs 20 (Figure 9 indeed shows a strong correlation between CD and visibility) or to inaccuracies in the dataset’s camera calibration process.
@@ 6.2.2 Results on LUCES-MV
Figure 9: Correlation between CD and visibility. From left to right: CD (in mm), number of cameras observing each vertex for LUCES-MV’s Bowl object, and CD vs visibility graph over the entire dataset.
@@ 6.2.3 Results on Skoltech3D
As for LUCES-MV, we compared our results on Skoltech3D against SuperNormal, which represents the state-of-the-art in MVPS. The CD reported in Table 5 emphasize the significant superiority of our results, as well as the interest of accounting for reflectance. Nevertheless, the results on this dataset are globally disappointing, as can be seen in Figure 10 on two examples exhibiting severe breaks in the surface.
@@ Methods
This is confirmed quantitatively by observing the strikingly high CD values. With similar spatial resolution and number of views as DiLiGenT-MV, the CD scores are drastically worse: the best-performing baseline’s CD ($1.332$ mm) being around 6×6$\times$6 × higher than for DiLiGenT-MV ($0.209$ mm). This time, we believe that it is the quality of the PS normals (in over-saturated areas, notably) that is the bottleneck – as the very high PS errors from Table 1 tend to indicate.
@@ 6.2.5 Sparse-View Scenario
One key advantage of using photometric stereo data is the richness of normal information. When these normals are of sufficiently high quality, as it is the case for the DiLiGenT-MV dataset, they impose strong constraints on the optimization process, enabling the proposed approach to perform well even in the sparse-view scenario, which is highly challenging for single-light MVS methods. This is assessed in Table 7, which reports the CD evolution when reducing the number of views from 20 to 5, for a panel of recent MVS methods and for the proposed one.
@@ 6.3 Ablations
Lastly, ablation studies were conducted on DiLiGenT-MV (Table 8) and LUCES-MV (Table 9) to quantify the impact of architectural choices within our method. More specifically, we compared the choice of $p$ in the NVR loss $\mathcal{L}_{\text{NVR}}^{p}$ (see Equation (8)), of considering reflectance maps (R) in addition to normal maps, of using the pixel-wise optimal lighting triplets or the canonical ones (O vs C), and of using reflectance embedding (+) instead of a standard reflectance parametrization.
Lighting optimization appears important, yet to a less extent than the choice of $p$: using $\mathcal{L}_{\text{NVR}}^{2}$ reduces CD scores by 9%. However, $\mathcal{L}_{\text{NVR}}^{1}$ performs better when normals and reflectance are used together, while reflectance not always contributes positively using $\mathcal{L}_{\text{NVR}}^{2}$. Finally, reflectance embedding (+) only slightly influences the metrics, which was to be expected since singular reflectance points are very sparse.
@@
$\mathcal{L}^{2}$+ 0.218 0.222 0.180 0.143 0.284 0.209
$\mathcal{L}^{2}$+ 0.214 0.219 0.179 0.153 0.291 0.211
$\mathcal{L}^{2}$ 0.219 0.230 0.176 0.174 0.304 0.221
$\mathcal{L}^{2}$ 0.156 0.219 0.187 0.134 0.276 0.194
$\mathcal{L}^{1}$+ 0.212 0.232 0.255 0.136 0.305 0.228
$\mathcal{L}^{1}$+ 0.196 0.235 0.271 0.164 0.291 0.231
$\mathcal{L}^{1}$ 0.214 0.231 0.305 0.140 0.314 0.241
$\mathcal{L}^{1}$ 0.194 0.234 0.351 0.160 0.324 0.252
@@
$\mathcal{L}^{2}$+ 0.645 0.623 0.203 0.583 0.309 0.309 0.518 0.292 0.260 0.279 0.402
$\mathcal{L}^{2}$+ 0.630 0.595 0.226 0.563 0.356 0.376 0.540 0.312 0.336 0.365 0.430
$\mathcal{L}^{2}$ 0.642 0.628 0.190 0.570 0.336 0.310 0.514 0.246 0.337 0.314 0.409
$\mathcal{L}^{2}$ 0.624 0.757 0.230 0.609 0.339 0.335 0.542 0.321 0.273 0.276 0.431
$\mathcal{L}^{1}$+ 0.764 0.834 0.264 0.757 0.295 0.384 0.524 0.333 0.308 0.287 0.475
$\mathcal{L}^{1}$+ 0.801 0.891 0.266 0.758 0.303 0.387 0.581 0.420 0.368 0.298 0.507
$\mathcal{L}^{1}$ 0.762 0.813 0.235 0.682 0.263 0.356 0.489 0.316 0.295 0.286 0.450
$\mathcal{L}^{1}$ 0.873 0.899 0.332 0.746 0.434 0.437 0.580 0.415 0.367 0.317 0.540

### v3

@@ 5.4 Reflectance Singularities
In comparison with SuperNormal, our approach has the advantage of taking into account not only normals, but also reflectance. However, this may sometimes turn into a drawback. Indeed, the actual reflectance values weight the optimisation process, potentially leading to singularities e.g., in the presence of very dark materials. Let us demonstrate this under the Lambertian assumption, in the case of grey scale images ($q=1$).
For simplicity, let us consider an ideal sampling scenario where the weight $w(\cdot,f(\mathbf{x}_{k}(\cdot)))$ in (9) is a Dirac function $\delta_{t_{k}}(\cdot)$, i.e. it is null everywhere except at some $t_{k}$ such that $\mathbf{x}_{k}:=\mathbf{x}_{k}(t_{k})$ lies on surface $\mathcal{S}$. By combining (3) and (10), the volumetric rendering (9) then simplifies to $\tilde{\mathbf{v}}_{k}(f,\boldsymbol{\rho})=\mathsf{L}_{k}\,\nabla f(\mathbf{x}_{k})\,\boldsymbol{\rho}(\mathbf{x}_{k})^{\top}$. If $q=1$, introducing the scalar variable $R$ such that $\boldsymbol{\rho}(\cdot)=\mathbf{r}\,R(\cdot)$, the NVR loss (8) can be rewritten:
$\mathcal{L}_{\text{NVR}}^{p}(f,R)=\sum_{k=1}^{m}|\mathbf{r}_{k}|\left\|R(\mathbf{x}_{k})\mathsf{L}_{k}\nabla f(\mathbf{x}_{k})-\mathsf{L}_{k}\,\mathbf{n}_{k}\right\|_{p}^{p}.$ (13)
This rewriting emphasises that reflectance acts as weighting factor in the loss function. Consequently, singular reflectance values may substantially influence the optimisation of geometry. In particular, dark colours essentially leave the geometry locally unconstrained, which may result in slow, sub-optimal convergence and a loss of fine details (see Figure 6-b).
@@
$\mathbf{r}_{k}^{+}$ $:=\frac{1}{q}\left[{\mathbf{r}_{k}}^{\top},~\left(q-\left\|\mathbf{r}_{k}\right\|_{p}^{p}\right)^{1/p}\right]^{\top}\ ,$ (14)
$\boldsymbol{\rho}^{+}(\cdot)$ $:=\frac{1}{q}\left[\boldsymbol{\rho}(\cdot)^{\top},~\left(q-\left\|\boldsymbol{\rho}(\cdot)\right\|_{p}^{p}\right)^{1/p}\right]^{\top}.$ (15)
It is easy to check that the norm of these vectors is equal to 1, whatever the values of $p$ and $q$. This reflectance embedding strategy prevents any unintended influence of singular reflectance values on geometry optimisation. This is qualitatively illustrated in Figure 6 on a synthetic experiment, using the $L_{1}$-norm ($p=1$) and grey level data ($q=1)$. Therein, the dark regions are not faithfully reconstructed in the standard case, whereas reflectance embedding significantly improves the results. Further experiments evaluating the interest of this embedding will be conducted in Section 6.
@@ 5.5 From RNb-NeuS to RNb-NeuS2
[…]
NeuS builds on NeRF [62] by introducing volumetric rendering with a transfer function that converts an SDF into a density function. This allows direct optimisation of scene geometry and extraction of a mesh via the SDF’s zero-level set. NeuS2 applies the same principles but replaces NeRF with Instant-NGP [64], which enables real-time rendering thanks to CUDA acceleration and the use of an optimisable hash grid for sampling during volumetric rendering. While CUDA contributes the most significant performance improvement, the hash grid further reduces computation time by focusing on surface-adjacent regions, enhancing detail preservation. NeuS2 retains these key optimisations while also including the SDF-based density representation from NeuS.
@@ 6.1.1 Evaluation Datasets
DiLiGenT-MV [53] is a benchmark of five real-world objects, some with complex surface profiles and reflectance. Each object is imaged from 20 calibrated viewpoints using the classical turntable MVPS acquisition setup [34], under 96 directional lighting conditions. Given the acquisition characteristics, the relatively low image resolution of $612\times 512$ pixels corresponds to approximately $0.4$ mm per pixel.
@@ 6.1.2 Photometric Stereo Methods
Among these methods, SDM-UniPS [38] (referred to as SDM in the result subsections) provides both normal and reflectance maps, enabling a complete evaluation of our pipeline. UniMS-PS [32] predicts slightly more accurate normals (see Table 1), yet no reflectance. It thus allows only for normal integration assessment, similar to SuperNormal [10]. To establish an upper bound on the achievable reconstruction quality, experiments were also conducted using ground truth normals.
In the forthcoming experiments, the notation N: X (resp. N/R: X) indicates that the inputs are normals alone (resp. normals and reflectance), estimated with method X.
Since architecture constraints limit SDM-UniPS to use only 10 PS images, a sampling strategy was employed. Specifically, each $\mathbf{n}_{k}$ and $\mathbf{r}_{k}$ was computed as the median of the estimated normals and reflectances across $100$ random trials, each trial involving 10 randomly chosen images. UniMS-PS scaling better thanks to multi-scale enhancement, sampling was not necessary.
@@ 6.2 MVPS Results
Methods Bear Buddha Cow Pot2 Reading Mean Approx. Time
Park16 0.932 0.375 0.335 0.972 0.525 0.628 N/A
Li20 $\dagger$ 0.206 0.257 0.101 0.213 0.255 0.206 N/A
Kaya22 0.380 0.396 0.297 0.390 0.344 0.361 N/A
PS-NeRF 0.253 0.304 0.275 0.248 0.353 0.287 $\sim$8-22h
Kaya23 0.320 0.202 0.209 0.370 0.270 0.274 N/A
MVPSNet 0.288 0.272 0.244 0.296 0.243 0.269 N/A
SuperNormal (N: SDM) 0.177 0.206 0.186 0.141 0.222 0.186 $\sim$5min
NLP-MVPS 0.211 0.176 0.169 0.213 0.250 0.204 N/A
RNb-NeuS (N/R: SDM) 0.255 0.216 0.282 0.175 0.276 0.241 $\sim$15h
Ours (N: CNN-PS) 0.160 0.197 0.242 0.166 0.283 0.210 $\sim$5min
Ours (N: SDPS-Net) 0.182 0.196 0.140 0.141 0.240 0.180
@@
Table 2: Chamfer distance (lower is better) and approximate training times on the DiLiGenT-MV dataset. Best results. Second best. Since $\dagger$ requires manual efforts [53], it is not ranked.
@@ 6.2.1 Results on DiLiGenT-MV
Table 2 presents the CD evaluation for the DiLiGenT-MV dataset. As a preliminary observation, let us remark the significant discrepancy (around $13\%$) between our results and those from the conference paper (RNb-NeuS), using the SDM-UniPS normal and reflectance maps. This discrepancy is due to the change of norm in the NVR loss (8), from $p=1$ to $p=2$, which severely impacts the CD scores. As shown in the ablation study (Table 8), using the $\mathcal{L}_{\text{NVR}}^{1}$ loss yields scores comparable with the conference paper.
Another notable observation is that whatever the PS method, our results are on par with state-of-the-art, or even better. Nevertheless, combining our method with UniMS-PS yields better results ($14\%$) than with SDM-UniPS, consistent with the higher accuracy of the former’s normals (see the first row of Table 1). This highlights the versatility of the proposed method, which adapts well to PS advancements.
@@
6.2.2 Results on LUCES-MV
Experiments on the LUCES-MV dataset (Table 4) confirmed our state-of-the-art results, and here reflectance visibly improves the reconstruction. Interestingly, the CD scores are significantly higher in comparison with DiLiGenT-MV, despite a better pixel resolution: for the Bowl object, even using GT normals, the CD is above $0.3$ mm. We believe this could be due both to the sparser number of views — 12 vs 20 (Figure 9 indeed shows a strong correlation between CD and visibility) or to inaccuracies in the dataset’s camera calibration process.
@@ 6.2.2 Results on LUCES-MV
Figure 9: Correlation between CD and visibility. From left to right: CD (in $m$), number of cameras observing each vertex for LUCES-MV’s Bowl object, and CD vs visibility graph over the entire dataset.
@@ 6.2.3 Results on Skoltech3D
As for LUCES-MV, we compared our results on Skoltech3D against SuperNormal, which represents the state-of-the-art in MVPS. The CD reported in Table 5 emphasise the significant superiority of our results, as well as the interest of accounting for reflectance. Nevertheless, the results on this dataset are globally disappointing, as can be seen in Figure 10 on two examples exhibiting severe breaks in the surface.
@@ Methods
This is confirmed quantitatively by observing the strikingly high CD values. With similar spatial resolution and number of views as DiLiGenT-MV, the CD scores are drastically worse: the best-performing baseline’s CD ($1.332$ mm) being around $6\times$ higher than for DiLiGenT-MV ($0.209$ mm). This time, we believe that it is the quality of the PS normals (in over-saturated areas, notably) that is the bottleneck – as the very high PS errors from Table 1 tend to indicate.
@@ 6.2.5 Sparse-View Scenario
One key advantage of using photometric stereo data is the richness of normal information. When these normals are of sufficiently high quality, as it is the case for the DiLiGenT-MV dataset, they impose strong constraints on the optimisation process, enabling the proposed approach to perform well even in the sparse-view scenario, which is highly challenging for single-light MVS methods. This is assessed in Table 7, which reports the CD evolution when reducing the number of views from 20 to 5, for a panel of recent MVS methods and for the proposed one.
@@ 6.3 Ablations
Lastly, ablation studies were conducted on DiLiGenT-MV (Table 8) and LUCES-MV (Table 9) to quantify the impact of architectural choices within our method. More specifically, we compared the choice of $p$ in the NVR loss $\mathcal{L}_{\text{NVR}}^{p}$ (Equation (8)), of considering reflectance maps (R) in addition to normal maps, of using the pixel-wise optimal lighting triplets or the canonical ones (O vs C), and of using reflectance embedding (+) instead of a standard reflectance parametrisation.
Lighting optimisation appears important, yet to a less extent than the choice of $p$: using $\mathcal{L}_{\text{NVR}}^{2}$ reduces CD scores by 9%. However, $\mathcal{L}_{\text{NVR}}^{1}$ performs better when normals and reflectance are used together, while reflectance not always contributes positively using $\mathcal{L}_{\text{NVR}}^{2}$. Finally, reflectance embedding (+) only slightly influences the metrics, which was to be expected since singular reflectance points are very sparse.
@@
$\mathcal{L}^{2}$OR+ 0.218 0.222 0.180 0.143 0.284 0.209
$\mathcal{L}^{2}$CR+ 0.214 0.219 0.179 0.153 0.291 0.211
$\mathcal{L}^{2}$OR 0.219 0.230 0.176 0.174 0.304 0.221
$\mathcal{L}^{2}$CR 0.236 0.221 0.173 0.163 0.312 0.221
$\mathcal{L}^{2}$O 0.156 0.219 0.187 0.134 0.276 0.194
$\mathcal{L}^{2}$C 0.166 0.219 0.181 0.137 0.276 0.196
$\mathcal{L}^{1}$OR+ 0.212 0.232 0.255 0.136 0.305 0.228
$\mathcal{L}^{1}$CR+ 0.196 0.235 0.271 0.164 0.291 0.231
$\mathcal{L}^{1}$OR 0.214 0.231 0.305 0.140 0.314 0.241
$\mathcal{L}^{1}$CR 0.185 0.237 0.301 0.166 0.315 0.241
$\mathcal{L}^{1}$O 0.194 0.234 0.351 0.160 0.324 0.252
$\mathcal{L}^{1}$C 0.183 0.238 0.388 0.202 0.326 0.267
@@
$\mathcal{L}^{2}$OR+ 0.645 0.623 0.203 0.583 0.309 0.309 0.518 0.292 0.260 0.279 0.402
$\mathcal{L}^{2}$CR+ 0.630 0.595 0.226 0.563 0.356 0.376 0.540 0.312 0.336 0.365 0.430
$\mathcal{L}^{2}$OR 0.642 0.628 0.190 0.570 0.336 0.310 0.514 0.246 0.337 0.314 0.409
$\mathcal{L}^{2}$CR 0.647 0.638 0.247 0.563 0.305 0.318 0.525 0.269 0.307 0.387 0.421
$\mathcal{L}^{2}$O 0.624 0.757 0.230 0.609 0.339 0.335 0.542 0.321 0.273 0.276 0.431
$\mathcal{L}^{2}$C 0.626 0.744 0.230 0.608 0.340 0.350 0.566 0.360 0.296 0.293 0.441
$\mathcal{L}^{1}$OR+ 0.764 0.834 0.264 0.757 0.295 0.384 0.524 0.333 0.308 0.287 0.475
$\mathcal{L}^{1}$CR+ 0.801 0.891 0.266 0.758 0.303 0.387 0.581 0.420 0.368 0.298 0.507
$\mathcal{L}^{1}$OR 0.762 0.813 0.235 0.682 0.263 0.356 0.489 0.316 0.295 0.286 0.450
$\mathcal{L}^{1}$CR 0.797 0.769 0.240 0.740 0.365 0.347 0.531 0.379 0.348 0.296 0.481
$\mathcal{L}^{1}$O 0.873 0.899 0.332 0.746 0.434 0.437 0.580 0.415 0.367 0.317 0.540
$\mathcal{L}^{1}$C 0.920 0.948 0.329 0.693 0.410 0.459 0.635 0.478 0.425 0.338 0.564